OpenStudy (anonymous):

A car travels 889 km in 15 days. At this rate, how far would it travel in 45 days?

OpenStudy (anonymous):

HMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMM c:

OpenStudy (anonymous):

c;

OpenStudy (aakashsudhakar):

45/15 = 3 889 * 3 = 2667 The car would travel 2667 km in 45 days if traveling at the same rate.

OpenStudy (texaschic101):

889/15 = x/45 (889 km in 15 days = x km in 45 days) cross multiply (15)(x) = (889)(45) 15x = 40005 -- divide both sides by 15 x = 40005/15 x = 2667 km. AaksahSudhakar is correct.....it is 2667 km in 45 days.